# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

### v1

@@ Appendix H Lipschitz process
$=\langle\nabla(\psi+g)(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle+\langle\nabla\psi(x_{\psi}),x_{\psi}-x_{\psi+g}\rangle-\langle\nabla g(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle$
@@
Lemma H.2 (Lipschitzness of minimizer of perturbed ERM).
@@
$C_{1}$ $=\left(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B\right),\qquad C_{2}=B+\beta_{\textnormal{max}}C_{1}.$ (10)
@@
First, consider $\hat{h}_{\beta}$. Fix $\beta>\beta^{\prime}$ in $\mathcal{I}$. Recalling the definition of $\hat{L}$ from (3), the difference between the objectives $\hat{L}(h,\beta)$ and $\hat{L}(h,\beta^{\prime})$ is the quadratic
@@
By Lemma N.3, the second term is convex, and since the third term is linear, it too is convex. The Hessian of the quadratic first term is $2\lambda(\beta^{\prime})^{2}\hat{\Sigma}$, from which it follows that $\psi$ is $2\lambda(\beta^{\prime})^{2}\lambda_{\textnormal{min}}(\hat{\Sigma})$-strongly convex.
Thus $\psi$ and $g$ satisfy the conditions of Lemma H.1, which implies the bound
@@
$\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(2\lambda\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}\left(\lambda(4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+2C_{\Phi})|\beta-\beta^{\prime}|+8\lambda\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B|\beta-\beta^{\prime}|\right)$
@@
$C_{1}=(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}\left((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B\right).$
@@
Lemma H.3 (Lipschitzness of minimizer of perturbed auxiliary ERM).
@@
The proof is almost identical to Lemma H.2. ∎
@@
Recall $\bar{L}$ from (16), $\hat{L}$ from (3), $\tilde{L}$ from (15), and $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ from Section 4. For any fixed $\beta\in\mathcal{I},$ define $\hat{h}_{\beta}$ as the minimizer of $h\mapsto\hat{L}(h,\beta)$ over $\mathcal{H}_{B}$, which exists under the conditions of Theorem 4.2 due to our argument checking the convexity of $h\mapsto\hat{L}(h,\beta)$ in Term (I) in the proof of Theorem 4.2.
Lemma H.4.
@@
$K_{1,\lambda}:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi})\lambda=:a_{1}\lambda,\qquad K_{2,\lambda}:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}})=:a_{2}.$
@@
$|\ell_{\alpha}(\hat{h}_{\beta},S)-\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)|\leqslant(1-\alpha)|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}|\leqslant(1-\alpha)C_{\Phi}\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(1-\alpha)C_{\Phi}C_{1}|\beta-\beta^{\prime}|,$
where in the first step we applied Lemma N.2, in the second step we used 1 to apply Lemma N.4, and in the third step we used Lemma H.2. Thus the first term in Equation 13 is bounded by $(1-\alpha)C_{\Phi}\mathbb{E}_{1}[C_{1}]|\beta-\beta^{\prime}|$, and the second term in Equation 13 is bounded by $(1-\alpha)C_{\Phi}\hat{\mathbb{E}}_{1}[C_{1}]|\beta-\beta^{\prime}|$. Summing, we deduce that
@@
The integrands of the first and third terms of Appendix H can be uniformly bounded as
@@
where in the first step we used difference of squares, in the second step we used 1 to apply Lemma N.4, in the third step we applied Lemma H.2 to bound the first factor and the triangle inequality and the bounds $\beta\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$ and $\|h\|\leqslant B$ for $h\in\mathcal{H}_{B}$ to bound the second factor. The integrand of the second and fourth term in (H) can be bounded as $|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|\leqslant C_{\Phi}\|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}\|\leqslant C_{\Phi}C_{2}|\beta-\beta^{\prime}|$, where in the first step we used 1 to apply Lemma N.4, and in the second step we applied Lemma H.2.
Plugging these into our bound in Appendix H, we deduce
@@
Appendix I Proof of Proposition 4.1
Fix $\lambda\geqslant 0$. Under the assumptions of Lemma L.3, there exists a global minimizer $(h^{*},\beta^{*})$ of $L(h,\beta)$. The first order condition with respect to $\beta$ reads $02\lambda\mathbb{E}_{1}[h^{*}(X)(\beta^{*}h^{*}(X)-r(X))]=$ 0. By Lemma N.5, the first order condition with respect to $h$ reads
@@
$=\mathbb{E}_{1}[h^{*}(X)(\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]+2\lambda\mathbb{E}_{1}[\beta^{*}(r_{\mathcal{H}}(X)-\beta^{*}h^{*}(X))(\beta^{*}h^{*}(X)-r(X))]$
@@
Appendix J Proof of Theorem 4.2
@@
We study the generalization error by conditioning on the unlabeled calibration or test data. Then our regularization becomes data-independent. Conditional on $\mathcal{S}_{1}$, Term (I) can be handled with Lemma G.1 above. Conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, Term (II) can be handled with Lemma G.1 above. Terms (III) and (IV) are empirical processes at fixed functions, conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$.
Term (I): We work conditional on $\mathcal{S}_{1}$. First, note that due to the definition of $\hat{L}$ from (3), we can write for any $(h,\beta)$,
@@
Since $\bar{L}(h,\beta)-\hat{L}(h,\beta)$ can be viewed as a difference of a population risk $\lambda\mathbb{E}_{3}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h]$ and an empirical risk $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$ with “regularizer” $\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]$, this expression enables us to apply Lemma G.1 to bound $\bar{L}(\hat{h},\hat{\beta})-\hat{L}(\hat{h},\hat{\beta})$.
@@
Hence, fixing $\beta$, we can apply Lemma G.1, choosing $m_{1}=n_{3}$ and $m_{2}=n_{2}$. Further, we choose ~$\tilde{\mathcal{H}}:=\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ with the norm $\langle\gamma,\Phi\rangle=\|\gamma\|_{2}$. Moreover, letting $z=(x^{\prime\prime},x^{\prime})$ for $x^{\prime\prime},x^{\prime}\in\mathcal{X}$, and $\xi=1/\lambda$, we use the objective function given by $(h,z)\mapsto f_{1}(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{J}(h,\beta,z)=\mathcal{J}_{1}(h,\beta,z)+\mathcal{J}_{2}(h,\beta,z)$, and where
@@
We now check the conditions of Lemma G.1.
Boundedness: Note that $|\mathcal{J}_{1}(h,\beta,z)|=|\beta|^{2}|h(x^{\prime\prime})|^{2}\leqslant\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4. Similarly, note that $|\mathcal{J}_{2}(h,\beta,z)|=2|\beta||h(x^{\prime})|\leqslant 2\beta_{\textnormal{max}}BC_{\Phi}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4. Thus $|\mathcal{J}_{1}(h,\beta,z)|$ and $|\mathcal{J}_{2}(h,\beta,z)|$ are both bounded by the sum $\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}+2\beta_{\textnormal{max}}BC_{\Phi}$.
@@
Strong convexity: Since $h\mapsto\ell_{\alpha}(h,s)$ is convex for all $s\in\mathbb{R}$ by Lemma N.3 and since $h\mapsto\hat{\mathbb{E}}_{2}[\beta h]$ is linear, the map $h\mapsto\xi\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]-2\hat{\mathbb{E}}_{2}[\beta h]$ is convex. Consider the map $h\mapsto\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]$. Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, this can be rewritten as $\gamma\mapsto\beta^{2}\gamma^{\top}\hat{\Sigma}\gamma$, a quadratic whose Hessian equals $2\beta^{2}\hat{\Sigma}$. By $\beta\geqslant\beta_{\textnormal{min}}$ for $\beta\in\mathcal{I}$ and 3, it follows that with probability $1-o(n_{3}^{-1})=1-o(n_{2}^{-1}+n_{3}^{-1})$, the map $h\mapsto\hat{\mathbb{E}}_{2,3}[f_{1}(h,Z)]$ is $\mu_{1}$-strongly convex, where $Z=(X^{\prime\prime},X^{\prime})$ with $X^{\prime}$ is uniform over $\mathcal{X}_{2}$ and $X^{\prime\prime}$ is uniform over $\mathcal{X}_{3}$, and where $\mu_{1}:=2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}$. In particular, $h\mapsto\frac{1}{\lambda}\hat{L}(h,\beta)$ is convex.
Let $\widetilde{C}_{1}=\frac{4\rho_{1}^{2}}{\mu_{1}}$. Let $K_{1}$ denote the Lipschitz constant of the process $\bar{W}_{\beta}$, where $K_{1}\leqslant K_{1,\lambda}$ with probability tending to unity conditional on $\mathcal{S}_{1}$ by 4 and Lemma H.4. From Lemma G.1 applied with $\xi=1/\lambda$, $\mathcal{L}=\frac{1}{\lambda}\bar{L}$, and ^$\hat{\mathcal{L}}=\frac{1}{\lambda}\hat{L}$, and $W=(\bar{L}-\hat{L})/\lambda$, we obtain that conditional on $\mathcal{S}_{1}$, for sufficiently large $n_{2},n_{3}$, with probability at least $1-\frac{\delta}{4}$, we have for Term (I) from (18),
@@
Term (II): We work conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$. The risks $\hat{L}$ and $\tilde{L}$ share the same data-independent regularization $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$. Write $z=(x,s)$ for $x\in\mathcal{X}$ and $s\in[0,1]$. Fixing $\beta$, we apply Lemma G.1 with the objective function $(h,z)\mapsto f(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where
@@
Since the empirical risk $\hat{L}$ is computed over the i.i.d. sample $Z_{i}=(X_{i},S_{i})$ for $i\in[n_{1}]$, we use the modified version of Lemma G.1 given in Remark 1. In particular, we check boundedness, convexity, and Lipschitzness of $\mathcal{J}$ without writing it as a sum $\mathcal{J}_{1}+\mathcal{J}_{2}$.
@@
where in the first step we used Lemma N.1, in the second step we used the triangle inequality and $s\in[0,1]$, and in the third step we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4.
Convexity: By Lemma N.3, $h\mapsto\mathcal{J}(h,\beta,z)$ is convex.
@@
$|\mathcal{J}(h,\beta,z)-\mathcal{J}(h,\beta,z)|=|\ell_{\alpha}(h(x),s)-\ell_{\alpha}(h^{\prime}(x),s)|\leqslant(1-\alpha)|h(x)-h^{\prime}(x)|\leqslant(1-\alpha)C_{\Phi}\|h-h^{\prime}\|,$
where in the second step we used Lemma N.2, and in the third step we used 1 to apply Lemma N.4. Thus $h\mapsto\mathcal{J}(h,\beta,z)$ is $\rho_{2}$-Lipschitz, where $\rho_{2}:=(1-\alpha)C_{\Phi}$.
@@
Let $\widetilde{C}_{2}(\lambda)=\frac{4\rho_{2}^{2}}{\mu_{2}(\lambda)}$. Let $K_{2}$ denote the Lipschitz constant of the process $W_{\beta}$; recall that conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, $K_{2}\leqslant K_{2,\lambda}$ deterministically on the event $C_{1}\leqslant C_{1,\textnormal{max}}$ by Lemma H.4. By the version of Lemma G.1 given in Remark 1, conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$, if $h\mapsto\mathcal{R}(h,\beta)$ is $\mu_{2}(\lambda)$-strongly convex, and if $C_{1}\leqslant C_{1,\textnormal{max}}$, then for sufficiently large $n_{1}$, with probability at least $1-\frac{\delta}{8}$, we have
@@
Since $\|h^{*}\|\leqslant B$ by $h^{*}\in\mathcal{H}_{B}$ and since 1 holds, we may apply Lemma N.4 to deduce that $\sup_{x\in\mathcal{X}}|h^{*}(x)|\leqslant BC_{\Phi}$. Consequently, for $\beta\in\mathcal{I}$, we have the uniform bound $\sup_{x\in\mathcal{X}}|\beta h^{*}(x)|\leqslant\beta_{\textnormal{max}}BC_{\Phi}$. By Hoeffding’s inequality (Hoeffding, , 1963), with probability at least $1-\frac{\delta}{8}$, we have
@@
Appendix K Proof of Theorem 4.3
@@
Lemma K.1 (Bounded suboptimality implies bounded gradient for smooth functions).
@@
By 1 and 5, we may apply Lemma N.5 to deduce that the Hessian of our population risk $L$ from (2) in the basis $\{\phi_{1},\ldots,\phi_{d}\}$ is the block matrix

### v2

@@ Appendix I Lipschitz process
$=\langle\nabla(\psi+g)(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle+\langle\nabla\psi(x_{\psi}),x_{\psi}-x_{\psi+g}\rangle$
$-\langle\nabla g(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle$
@@
Lemma I.2 (Lipschitzness of minimizer of perturbed ERM).
@@
$C_{1}$ $=(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B),\quad C_{2}=B+\beta_{\textnormal{max}}C_{1}.$ (8)
@@
First, consider $\hat{h}_{\beta}$. Fix $\beta>\beta^{\prime}$ in $\mathcal{I}$. Recalling the definition of $\hat{L}$ from (Empirical-LR-QR), the difference between the objectives $\hat{L}(h,\beta)$ and $\hat{L}(h,\beta^{\prime})$ is the quadratic
@@
By Lemma O.3, the second term is convex, and since the third term is linear, it too is convex. The Hessian of the quadratic first term is $2\lambda(\beta^{\prime})^{2}\hat{\Sigma}$, from which it follows that $\psi$ is $2\lambda(\beta^{\prime})^{2}\lambda_{\textnormal{min}}(\hat{\Sigma})$-strongly convex.
Thus $\psi$ and $g$ satisfy the conditions of Lemma I.1, which implies the bound
@@
$\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|$$\leqslant$
$(2\lambda\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}(\lambda(4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+2C_{\Phi})|\beta-\beta^{\prime}|+8\lambda\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B|\beta-\beta^{\prime}|).$
@@
$C_{1}=(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B).$
@@
Lemma I.3 (Lipschitzness of minimizer of perturbed auxiliary ERM).
@@
The proof is almost identical to Lemma I.2. ∎
@@
Recall $\bar{L}$ from (14), $\hat{L}$ from (Empirical-LR-QR), $\tilde{L}$ from (13), and $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ from Section 4. For any fixed $\beta\in\mathcal{I},$ define $\hat{h}_{\beta}$ as the minimizer of $h\mapsto\hat{L}(h,\beta)$ over $\mathcal{H}_{B}$, which exists under the conditions of Theorem 4.2 due to our argument checking the convexity of $h\mapsto\hat{L}(h,\beta)$ in Term (I) in the proof of Theorem 4.2.
Lemma I.4.
@@
$K_{1,\lambda}$ $:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi})\lambda=:a_{1}\lambda,$
$K_{2,\lambda}$ $:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}})=:a_{2},$
@@
$|\ell_{\alpha}(\hat{h}_{\beta},S)-\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)|$ $\leqslant(1-\alpha)|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}|$
$\leqslant(1-\alpha)C_{\Phi}\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(1-\alpha)C_{\Phi}C_{1}|\beta-\beta^{\prime}|,$
where in the first step we applied Lemma O.2, in the second step we used 1 to apply Lemma O.4, and in the third step we used Lemma I.2. Thus the first term in Equation 11 is bounded by $(1-\alpha)C_{\Phi}\mathbb{E}_{1}[C_{1}]|\beta-\beta^{\prime}|$, and the second term in Equation 11 is bounded by $(1-\alpha)C_{\Phi}\hat{\mathbb{E}}_{1}[C_{1}]|\beta-\beta^{\prime}|$. Summing, we deduce that
@@
The integrands of the first and third terms of Appendix I can be uniformly bounded as
@@
where in the first step we used difference of squares, in the second step we used 1 to apply Lemma O.4, in the third step we applied Lemma I.2 to bound the first factor and the triangle inequality and the bounds $\beta\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$ and $\|h\|\leqslant B$ for $h\in\mathcal{H}_{B}$ to bound the second factor. The integrand of the second and fourth term in (I) can be bounded as $|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|\leqslant C_{\Phi}\|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}\|\leqslant C_{\Phi}C_{2}|\beta-\beta^{\prime}|$, where in the first step we used 1 to apply Lemma O.4, and in the second step we applied Lemma I.2.
Plugging these into our bound in Appendix I, we deduce
@@
Appendix J Proof of Proposition 4.1
Fix $\lambda\geqslant 0$. Under the assumptions of Lemma M.3, there exists a global minimizer $(h^{*},\beta^{*})$ of $L(h,\beta)$. The first order condition with respect to $\beta$ reads $02\lambda\mathbb{E}_{1}[h^{*}(X)(\beta^{*}h^{*}(X)-r(X))]=$ 0. By Lemma O.5, the first order condition with respect to $h$ reads
@@
$=\mathbb{E}_{1}[h^{*}(X)(\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]$
$+2\lambda\mathbb{E}_{1}[\beta^{*}(r_{\mathcal{H}}(X)-\beta^{*}h^{*}(X))(\beta^{*}h^{*}(X)-r(X))]$
@@
Appendix K Proof of Theorem 4.2
@@
We study the generalization error by conditioning on the unlabeled calibration or test data. Then our regularization becomes data-independent. Conditional on $\mathcal{S}_{1}$, Term (I) can be handled with Lemma H.1 above. Conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, Term (II) can be handled with Lemma H.1 above. Terms (III) and (IV) are empirical processes at fixed functions, conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$.
Term (I): We work conditional on $\mathcal{S}_{1}$. First, note that due to the definition of $\hat{L}$ from (Empirical-LR-QR), we can write for any $(h,\beta)$,
@@
Since $\bar{L}(h,\beta)-\hat{L}(h,\beta)$ can be viewed as a difference of a population risk $\lambda\mathbb{E}_{3}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h]$ and an empirical risk $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$ with “regularizer” $\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]$, this expression enables us to apply Lemma H.1 to bound $\bar{L}(\hat{h},\hat{\beta})-\hat{L}(\hat{h},\hat{\beta})$.
@@
Hence, fixing $\beta$, we can apply Lemma H.1, choosing $m_{1}=n_{3}$ and $m_{2}=n_{2}$. Further, we choose ~$\tilde{\mathcal{H}}:=\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ with the norm $\langle\gamma,\Phi\rangle=\|\gamma\|_{2}$. Moreover, letting $z=(x^{\prime\prime},x^{\prime})$ for $x^{\prime\prime},x^{\prime}\in\mathcal{X}$, and $\xi=1/\lambda$, we use the objective function given by $(h,z)\mapsto f_{1}(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{J}(h,\beta,z)=\mathcal{J}_{1}(h,\beta,z)+\mathcal{J}_{2}(h,\beta,z)$, and where
@@
We now check the conditions of Lemma H.1.
Boundedness: Note that $|\mathcal{J}_{1}(h,\beta,z)|=|\beta|^{2}|h(x^{\prime\prime})|^{2}\leqslant\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma O.4. Similarly, note that $|\mathcal{J}_{2}(h,\beta,z)|=2|\beta||h(x^{\prime})|\leqslant 2\beta_{\textnormal{max}}BC_{\Phi}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma O.4. Thus $|\mathcal{J}_{1}(h,\beta,z)|$ and $|\mathcal{J}_{2}(h,\beta,z)|$ are both bounded by the sum $\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}+2\beta_{\textnormal{max}}BC_{\Phi}$.
@@
Strong convexity: Since $h\mapsto\ell_{\alpha}(h,s)$ is convex for all $s\in\mathbb{R}$ by Lemma O.3 and since $h\mapsto\hat{\mathbb{E}}_{2}[\beta h]$ is linear, the map $h\mapsto\xi\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]-2\hat{\mathbb{E}}_{2}[\beta h]$ is convex. Consider the map $h\mapsto\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]$. Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, this can be rewritten as $\gamma\mapsto\beta^{2}\gamma^{\top}\hat{\Sigma}\gamma$, a quadratic whose Hessian equals $2\beta^{2}\hat{\Sigma}$. By $\beta\geqslant\beta_{\textnormal{min}}$ for $\beta\in\mathcal{I}$ and 3, it follows that with probability $1-o(n_{3}^{-1})=1-o(n_{2}^{-1}+n_{3}^{-1})$, the map $h\mapsto\hat{\mathbb{E}}_{2,3}[f_{1}(h,Z)]$ is $\mu_{1}$-strongly convex, where $Z=(X^{\prime\prime},X^{\prime})$ with $X^{\prime}$ is uniform over $\mathcal{X}_{2}$ and $X^{\prime\prime}$ is uniform over $\mathcal{X}_{3}$, and where $\mu_{1}:=2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}$. In particular, $h\mapsto\frac{1}{\lambda}\hat{L}(h,\beta)$ is convex.
Let $\widetilde{C}_{1}=\frac{4\rho_{1}^{2}}{\mu_{1}}$. Let $K_{1}$ denote the Lipschitz constant of the process $\bar{W}_{\beta}$, where $K_{1}\leqslant K_{1,\lambda}$ with probability tending to unity conditional on $\mathcal{S}_{1}$ by 4 and Lemma I.4. From Lemma H.1 applied with $\xi=1/\lambda$, $\mathcal{L}=\frac{1}{\lambda}\bar{L}$, and ^$\hat{\mathcal{L}}=\frac{1}{\lambda}\hat{L}$, and $W=(\bar{L}-\hat{L})/\lambda$, we obtain that conditional on $\mathcal{S}_{1}$, for sufficiently large $n_{2},n_{3}$, with probability at least $1-\frac{\delta}{4}$, we have for Term (I) from (16),
@@
Term (II): We work conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$. The risks $\hat{L}$ and $\tilde{L}$ share the same data-independent regularization $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$. Write $z=(x,s)$ for $x\in\mathcal{X}$ and $s\in[0,1]$. Fixing $\beta$, we apply Lemma H.1 with the objective function $(h,z)\mapsto f(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where
@@
Since the empirical risk $\hat{L}$ is computed over the i.i.d. sample $Z_{i}=(X_{i},S_{i})$ for $i\in[n_{1}]$, we use the modified version of Lemma H.1 given in Remark 1. In particular, we check boundedness, convexity, and Lipschitzness of $\mathcal{J}$ without writing it as a sum $\mathcal{J}_{1}+\mathcal{J}_{2}$.
@@
where in the first step we used Lemma O.1, in the second step we used the triangle inequality and $s\in[0,1]$, and in the third step we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma O.4.
Convexity: By Lemma O.3, $h\mapsto\mathcal{J}(h,\beta,z)$ is convex.
@@
$|\mathcal{J}(h,\beta,z)-\mathcal{J}(h,\beta,z)|=|\ell_{\alpha}(h(x),s)-\ell_{\alpha}(h^{\prime}(x),s)|$
$\leqslant(1-\alpha)|h(x)-h^{\prime}(x)|\leqslant(1-\alpha)C_{\Phi}\|h-h^{\prime}\|,$
where in the second step we used Lemma O.2, and in the third step we used 1 to apply Lemma O.4. Thus $h\mapsto\mathcal{J}(h,\beta,z)$ is $\rho_{2}$-Lipschitz, where $\rho_{2}:=(1-\alpha)C_{\Phi}$.
@@
Let $\widetilde{C}_{2}(\lambda)=\frac{4\rho_{2}^{2}}{\mu_{2}(\lambda)}$. Let $K_{2}$ denote the Lipschitz constant of the process $W_{\beta}$; recall that conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, $K_{2}\leqslant K_{2,\lambda}$ deterministically on the event $C_{1}\leqslant C_{1,\textnormal{max}}$ by Lemma I.4. By the version of Lemma H.1 given in Remark 1, conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$, if $h\mapsto\mathcal{R}(h,\beta)$ is $\mu_{2}(\lambda)$-strongly convex, and if $C_{1}\leqslant C_{1,\textnormal{max}}$, then for sufficiently large $n_{1}$, with probability at least $1-\frac{\delta}{8}$, we have
@@
Since $\|h^{*}\|\leqslant B$ by $h^{*}\in\mathcal{H}_{B}$ and since 1 holds, we may apply Lemma O.4 to deduce that $\sup_{x\in\mathcal{X}}|h^{*}(x)|\leqslant BC_{\Phi}$. Consequently, for $\beta\in\mathcal{I}$, we have the uniform bound $\sup_{x\in\mathcal{X}}|\beta h^{*}(x)|\leqslant\beta_{\textnormal{max}}BC_{\Phi}$. By Hoeffding’s inequality (Hoeffding, , 1963), with probability at least $1-\frac{\delta}{8}$, we have
@@
Appendix L Proof of Theorem 4.3
@@
Lemma L.1 (Bounded suboptimality implies bounded gradient for smooth functions).
@@
By 1 and 5, we may apply Lemma O.5 to deduce that the Hessian of our population risk $L$ from (LR-QR) in the basis $\{\phi_{1},\ldots,\phi_{d}\}$ is the block matrix